OpenStudy (anonymous):

If an investment of $2500 grows to $3255.65 in six years with interest compounded annually, what is the interest rate? (Round your answer to one decimal place.)

OpenStudy (anonymous):

Ive done probs like this except they give me the rate and dont give me the final $... How would I do this?

hero (hero):

You're still stuck on these?

OpenStudy (anonymous):

this ones different lol

OpenStudy (anonymous):

i get that you have to plug the #s in the formula but this one doesnt give you the rate from the get go

OpenStudy (anonymous):

p(1+r/n)^nt

OpenStudy (anonymous):

you there?

hero (hero):

I am here

OpenStudy (anonymous):

can u help?

hero (hero):

You have to find the rate

hero (hero):

Solve the given equation for r first. Then plug in everything

hero (hero):

Then simplify

OpenStudy (anonymous):

there is no given equation

hero (hero):

Yes there is. Says "compounded annually". There's a formula for that

OpenStudy (anonymous):

i know. i posted it above.

OpenStudy (anonymous):

could u start me off

OpenStudy (anonymous):

at least

OpenStudy (anonymous):

if an investment of $2500 grows to $3255.65 in six years with interest compounded annually, what is the interest rate? (Round your answer to one decimal place.)

OpenStudy (anonymous):

solve \[3255.56=2500(1+r)^6\] for r

OpenStudy (anonymous):

divide by 2500, take the sixth root, subtract 1. need a calculator

hero (hero):

t = 6, not 12

OpenStudy (anonymous):

i get 1.0449 so interest is \[r=0.0449=4.49\%\]

OpenStudy (anonymous):

hmm i didnt get that

OpenStudy (anonymous):

i got 3.87...

OpenStudy (anonymous):

i divivided3255.56 by 2500 and got 1.302224 then took the 6th root and got 4.876566723 and then subtracted 1 and got 3.876566723

OpenStudy (anonymous):

what am i doing wrong here

OpenStudy (anonymous):

hero? satellite?

OpenStudy (anonymous):

3255.56=2500(1+r)^6 First dived through by 2500 3255.56/2500=(1+r)^6 3255.56/2500=1.302224 So 1.302224=(1+r)^6 You have to take the sixth root now: \sqrt[6]{1.302224?}=\[1.0449952^6=1.302224\ because 1.0449952^6=1.302224 1.0449952=1+r so r=0.0449952 or 4.5% round up a tiny bit

OpenStudy (anonymous):

Sorry that equation did not want to show properly \[\sqrt[6]{1.302224?}=1.0449952^6=1.302224\]

OpenStudy (anonymous):

Anyway the sixth root is =1.045

OpenStudy (anonymous):

Sorry had a few beers with my friends :D

OpenStudy (anonymous):

haha i want one!

OpenStudy (anonymous):

my flatmate is getting some more I guess

OpenStudy (anonymous):

maybe not, he is gone for some time now....

OpenStudy (anonymous):

haha he's drinking it all himself

OpenStudy (anonymous):

might be, found a nice girl and could not leave the shop :)

OpenStudy (anonymous):

haha uh oh